OpenStudy (anonymous):

Keisha's parents want to save twenty thousand dollars in her college savings account over the next fifteen years. They have eight thousand dollars to use as an initial deposit. What simple annual interest rate do they need to meet their goal? Round your answer to the nearest tenth. two point five percent two point seven percent ten percent twelve percent

OpenStudy (texaschic101):

I = PRT looking for R, change the formula to I/PT = R P = 8000 T = 15 I = (20,000 - 8000) = 12000 now lets sub... I/PT = R 12,000/(8000)(15) = R 12,000/120,000 = R .1 = R (change to percent by moving decimal 2 spaces to the right) 10% = R